OpenStudy (anonymous):

voltage is given by v=100 sin(200πt+π/4) Use differential calculus, calculate the time taken for the voltage to reach its fist maximum.

OpenStudy (anonymous):

Since you're looking for a maximum, you're going to take the first derivative and then set it equal to zero. Then solve for t in that situation. \[dV=100(200\pi)\cos(200\pi t + \frac{\pi}{4})\] Note that the 200pi is coming from the chain rule. Now set it equal to zero and solve for t: \[0=100(200\pi)\cos(200\pi t + \frac{\pi}{4})\]

OpenStudy (anonymous):

\(dV/dt\)*

OpenStudy (anonymous):

Oops, right! Thanks for catching that.

OpenStudy (anonymous):

thank you very much for your help.